# Online Matroid Embeddings

### v1

@@ 3 Online Matroid Embeddings
Given a matroid with ordered ground set specified by a pair $(\textbf{M},\pi)$, we say that $(\textbf{M}^{\prime},\pi^{\prime})$ is a prefix-restriction of $(\textbf{M},\pi)$ if $n^{\prime}=|\textbf{M}^{\prime}|<|\textbf{M}|$, $\textbf{M}^{\prime}$ is the restriction of M to $\pi([n^{\prime}])$ and $\pi^{\prime}$ is the restriction of $\pi$ to $[n^{\prime}]$.
@@
for every $\textbf{M}\in\mathcal{C}$ and every indexing $\pi:[n]\rightarrow\textbf{M}$ of the ground set of M, such for every prefix-restriction $(\textbf{M}^{\prime},\pi^{\prime})$ of $(\textbf{M},\pi)$, the map $f_{\textbf{M}^{\prime},\pi^{\prime}}$ is the restriction of $f_{\textbf{M},\pi}$ to the ground set of $\textbf{M}^{\prime}$.
@@
We define $f^{\prime}_{\textbf{M},\pi}$ as follows: for each $u\in\textbf{M}$ if $u=\pi(k)$ let $f^{\prime}_{\textbf{M},\pi}(u)=(f_{\textbf{M},\pi}(u),j)$ where $j=|i\in[k];f_{\textbf{M},\pi}(\pi(i))=u|$. The functions $f^{\prime}_{\textbf{M},\pi}$ are injective by construction and they are matroid morphisms by the definition of $\textbf{BigM}_{[n]}$. In fact: $f_{\textbf{M},\pi}=\phi\circ f^{\prime}_{\textbf{M},\pi}$ where $\phi$ is the natural projection $_{[n]}\rightarrow\textbf{BigM}$ BigM. Finally note that they can be constructed online since the identity of the copy used is only a function of the set of elements that arrived up to this point. ∎
@@ 4 Binary Matroids
and hence we can map: $f(a)$ to $f(u_{1})+\ldots+f(u_{m})$ (recall that $1=-1$ in $\mathbb{F}_{2}$).
@@
[…]
@@
In this section we repeatedly use the fact that $\mathbb{F}_{2}$ has only one non-zero constant, so whenever we identify a circuit in the matroid, we know exactly what is the linear dependency between the elements in the corresponding vector field. This is no longer true even in slightly larger fields like $\mathbb{F}_{3}$. If vectors $u,v,w\in\mathbb{F}_{3}^{n}$ form a circuit, it could be that: $w=\pm u\pm v$ in the representation. As a consequence, given two matroid morphisms $f,g:\textbf{M}\rightarrow\mathbb{F}_{3}^{n}$ there may not exist a vector-space automorphism $A$ of $\mathbb{F}_{3}^{n}$ such that $f=Ag$. In the previous example, if $f(u)=g(u)$, $f(v)=g(v)$ but $f(w)=f(u)+f(v)$ but $g(w)=g(u)-g(v)$ no such automorphism can exist.
@@ 5 OMEs and the Matroid Secretary Problem
We consider two versions of the matroid secretary problem (MSP):
  • Known-matroid MSP: In the first version, a matroid M is known to the algorithm ahead of time. The only information missing is the weight of each element. The algorithm then processes pairs $(u,w_{u})$ in random order, where $u$ is an element in the ground set of M, and $w_{u}\in\mathbb{R}_{+}$ is its weight. Upon seeing the element, the algorithm must irrevocably decide whether to accept that element or not, subject to the constraint that the set of accepted elements must be an independent set of M.
  • Online-revealed-matroid MSP: In the second version, the algorithm has only access to the number of elements $n=|\textbf{M}|$ and to a promise that $\textbf{M}\in\mathcal{C}$ for a class of matroids $\mathcal{C}$. The algorithm again processes pairs $(u,w_{u})$ in random order at each time, but it only knows the rank function restricted to the subset of elements that have already arrived.
The first of these is identified as the RO-AA-MK model in the nomenclature of [29], whereas the second is identified as the RO-AA-MN model. The second version is clearly harder in the sense that an $\alpha$-approximation to the second version immediately implies an $\alpha$-approximation to the first version. Our goal is to understand when is the converse true. We will show the following theorem:
@@
For binary matroids, there is no gap in approximability of the known-matroid MSP and the online-revealed-matroid MSP.
The rest of this section is dedicated to a reduction from the known-matroid MSP for $\textbf{BigM}_{[n]}$ together with an online matroid embedding to the online-revealed-matroid MSP. The main challenge of the reduction is to (mostly) preserve random order of arrivals.
@@
Observe that if we ignore the weights, the order of the elements of BigM is the same in both processes since they are determined by the timestamps $t_{vj}$. What we are left to argue is that for each $v$ we select uniformly random timestamp $t_{vj}$ to assign the non-negative weight. For that, observe that since $f_{\textbf{M},\pi}$ is order independent, it has the same distribution as if we first sampled a monomorphism $f:\textbf{M}\rightarrow\textbf{BigM}$, and then we sampled an independent uniform indexing $\pi$ for the arrival order of the elements in $f(\textbf{M})$. This implies that when we assign the timestamps $T_{1},\dots,T_{n}$ according to $\pi$, the resulting distribution is the same as if we assigned i.i.d. Uniform$[0,1]$ timestamps $T_{v}$ to each element $v\in f(\textbf{M})$, and therefore, the interval $T_{v}$ lands in is uniformly chosen and independent across elements $v\in f(\textbf{M})$. Now fix a certain timestamp $t_{vj}$ and let $I_{i}$ be the interval containing it. We will show that the probability that this timestamp is selected is exactly $1/k$.
Consider two cases: either $X_{vi}\leq k/d$ in which case the probability of sampling $t_{vj}$ is the probability that the timestamp $T_{v}$ is in $I_{i}$ times (which is $1/d$), times the probability we decide to sample a timestamp inside $I_{i}$ (which is $X_{vi}d/k$), times the probability that out of those, we choose $t_{vj}$ (which is $1/X_{vi}$). The total probability is:
@@
Taking those lemmas together, we can conclude the proof of Theorem 5.1. For that, let Alg be an $\alpha$-competitive algorithm for $\textbf{BigM}_{[k]}$ and let $Y$ represent the sequence of the MSP sampled by the offline reduction. Let’s represent by $\textsf{Alg}(Y)$ the weight of the elements selected by Alg and Opt the weight of the optimal basis. By the fact that the offline reduction produces an instance in random order, we know that $\mathbb{E}[\textsf{Alg}(Y)]\geq\alpha\textsf{Opt}$.
@@
6 Graphic and Regular Matroids

### v2

@@ 3 Online Matroid Embeddings
Given a matroid with ordered ground set specified by a pair $(\textbf{M},\pi)$, we say that $(\textbf{M}^{\prime},\pi^{\prime})$ is a prefix-restriction of $(\textbf{M},\pi)$ if $n^{\prime}=|\textbf{M}^{\prime}|<|\textbf{M}|$, $\textbf{M}^{\prime}$ is the restriction of M to $\pi([n^{\prime}])$ and $\pi^{\prime}$ is the restriction of $\pi$ to $n^{\prime}]$.
@@
for every $\textbf{M}\in\mathcal{C}$ and every indexing $\pi:[n]\rightarrow\textbf{M}$ of the ground set of M, such that for every prefix-restriction $(\textbf{M}^{\prime},\pi^{\prime})$ of $(\textbf{M},\pi)$, the map $f_{\textbf{M}^{\prime},\pi^{\prime}}$ is the restriction of $f_{\textbf{M},\pi}$ to the ground set of $\textbf{M}^{\prime}$.
@@
We define $f^{\prime}_{\textbf{M},\pi}$ as follows: for each $u\in\textbf{M}$ if $u=\pi(k)$ let $f^{\prime}_{\textbf{M},\pi}(u)=(f_{\textbf{M},\pi}(u),j)$ where $j=|i\in[k];f_{\textbf{M},\pi}(\pi(i))=u|$. The functions $f^{\prime}_{\textbf{M},\pi}$ are injective by construction and they are matroid morphisms by the definition of $\textbf{BigM}_{[n]}$. In fact: $f_{\textbf{M},\pi}=\phi\circ f^{\prime}_{\textbf{M},\pi}$ where $\phi$ is the natural projection $\textbf{BigM}_{[n]}\rightarrow\textbf{BigM}$. Finally note that they can be constructed online since the identity of the copy used is only a function of the set of elements that arrived up to this point. ∎
@@
Uniform Order-Independent Embedding
Let $f$ be an order-independent online embedding from $\textbf{M}\rightarrow\textbf{BigM}$. We consider an order independent-randomized embedding $g:\textbf{M}\rightarrow\textbf{BigM}$ by composing $f$ with uniformly random automorphism $f^{\prime}\in\operatorname{Aut}(\textbf{BigM})$, i.e. $g=f^{\prime}\circ f$.
Interestingly, whenever BigM satisfies $2$-transitive property then the embedding $g$ maps each element of $e\in\textbf{M}$ uniformly at random over the matroid BigM. In addition, for any independent set of pair of elements $\{e,e^{\prime}\}\subseteq\textbf{M}$ and pair of elements $\{\tilde{e},\tilde{e}^{\prime}\}\subseteq\textbf{BigM}$, the events $\{g(e)=\tilde{e}\}$ and $\{g(e^{\prime})=\tilde{e}^{\prime}\}$ are ‘almost’ independent. More formally,
Lemma 3.2.
Given a simple host matroid (loop-free and no parallel elements) BigM such that for any two pairs of distinct elements $\{e_{1},e_{2}\}$ and $\{e_{1}^{\prime},e_{2}^{\prime}\}$ there exists an automorphism $f^{\prime}\in\operatorname{Aut}(\textbf{BigM})$ satisfying $f^{\prime}(e_{1})=e_{1}^{\prime}$ and $f^{\prime}(e_{2})=e_{2}^{\prime}$, then a uniformly random automorphism $f$ sampled from $\operatorname{Aut}(\textbf{BigM})$ satisfies:
  1. 1.
    For any $e,e^{\prime}\in\textbf{BigM}$, $\Pr[f(e)=e^{\prime}]=\frac{1}{n}$.
  2. 2.
    For any two pairs of elements $e_{1},e_{2}$ and $e_{1}^{\prime},e_{2}^{\prime}$ (s.t. $e_{1}\neq e_{2}$ and $e_{1}^{\prime}\neq e_{2}^{\prime}$), we have
    $\Pr[f(e_{1})=e_{1}^{\prime}\land f(e_{2})=e_{2}^{\prime}]=\frac{1}{n\cdot(n-1)}$
Proof.
Consider the action of the group of automorphisms $\operatorname{Aut}(\textbf{BigM})$ on the set of pairs of distinct elements $P=\{(e_{1},e_{2}):e_{1},e_{2}\in\textbf{BigM},e_{1}\neq e_{2}\}$. For two pairs $(e_{1},e_{2}),(e_{1}^{\prime},e_{2}^{\prime})\in P$, the set of automorphisms $f\in\operatorname{Aut}(\textbf{BigM})$ that satisfy $f(e_{1},e_{2})=(e_{1}^{\prime},e_{2}^{\prime})$ is nonempty, and therefore, it is a coset of the subgroup of stabilizers of $(e_{1},e_{2})$ (i.e., the set of automorphisms that satisfy $f(e_{1},e_{2})=(e_{1},e_{2})$). Since all cosets of a subgroup must have the same size, and because all pairs $(e_{1}^{\prime},e_{2}^{\prime})\in P$ define a different coset, for a uniformly drawn automorphism $f$,
$\Pr[f(e_{1},e_{2})=(e_{1}^{\prime},e_{2}^{\prime})]=\frac{1}{|P|}=\frac{1}{n\cdot(n-1)}.$
An analogous argument gives that $\Pr[f(e)=e^{\prime}]=1/n$. ∎
@@ 4 OMEs for Binary Matroids
and hence we can map: $f(a)$ to $f(u_{1})+\ldots+f(u_{m})$ (recall that 1=11=-1 in $\mathbb{F}_{2}$).
@@
Theorem 4.4 is not particular to binary matroids. The only fact it uses is that all the morphisms of the OMM belong to the same orbit under the action of the automorphism group $\text{Aut}(\textbf{BigM})$. We can state it more generally as follows. The proof is identical to Theorem 4.4, so we omit it here.
@@
In this section we repeatedly use the fact that $\mathbb{F}_{2}$ has only one non-zero constant, so whenever we identify a circuit in the matroid, we know exactly what is the linear dependency between the elements in the corresponding vector field. This is no longer true even in slightly larger fields like $\mathbb{F}_{3}$. If vectors $u,v,w\in\mathbb{F}_{3}^{n}$ form a circuit, it could be that: $w=\pm u\pm v$ in the representation. As a consequence, given two matroid morphisms $f,g:\textbf{M}\rightarrow\mathbb{F}_{3}^{n}$ there may not exist a vector-space automorphism $A$ of $\mathbb{F}_{3}^{n}$ such that $f=Ag$. In the previous example, if $f(u)=g(u)$, $f(v)=g(v)$ but $f(w)=f(u)+f(v)$ but g(w)=g(u)g(v)g(w)=g(u)-g(v) no such automorphism can exist.
@@ 5 OMEs and the Matroid Secretary Problem
In this section, we use OMEs to explore the complexity of the matroid secretary problem (MSP) on binary matroids. We consider three versions of the problem, each making different assumptions on the data generation process and what’s known to the algorithm. We will use OMEs to establish equivalences between these problems, showing that seemingly simpler problems are actually equivalent to harder ones (up to constants).
Three Versions of the MSP.
We consider the following three versions of the MSP, and aim to establish the relations in Figure 1. In all three variants, the goal is an algorithm for selecting elements that form an independent set, and whose combined weight is in expectation an $\alpha$-approximation to the weight of the optimal basis.
  • Online-revealed-matroid MSP: In this version of the problem, there is an underlying matroid M, which is a priori unknown to the algorithm. The algorithm has only access to the number of elements $n=|\textbf{M}|$ and to a promise that $\textbf{M}\in\mathcal{C}$ for a class of matroids $\mathcal{C}$. For each element $u\in\textbf{M}$, an adversary determines a weight $w_{u}\in\mathbb{R}_{+}$. The algorithm processes pairs $(u,w_{u})$ in random order at each time, but it only knows the rank function restricted to the subset of elements that have already arrived. Upon seeing the element, the algorithm must irrevocably decide whether to accept that element or not, subject to the constraint that the set of accepted elements must be an independent set of M.
  • Known-matroid MSP: In this version, the matroid M is known to the algorithm ahead of time. The only information missing is the weight of each element, which is again chosen adversarially. As before elements arrive in random order, and the algorithm must make immediate accept/reject decisions, with the restriction that the chosen set of elements must be an independent set of M.
  • Prophet MSP: In this version, the matroid M is again known to the algorithm ahead of time. However, this time the weight of each element is drawn from a known distribution $\mathcal{D}$, which can potentially sample weights in a correlated manner. As in the other versions elements are then presented to the algorithm in random order, and the algorithm aims to select an independent set of high weight in an online manner.
Note that the first version is clearly harder than the second and the second version is clearly harder than the third, in the sense that an $\alpha$-approximation to the harder problem immediately implies an $\alpha$-approximation to the simpler one. We derive approximate “inverses” of these comparisons from the existence of OMEs, even if we restrict prophet MSP to pairwise-independent distributions.
Our Reductions.
We first use OMEs to show an (essentially exact) “inverse” of the comparison between the online-revealed-matroid MSP and the known-matroid MSP, implying that for binary matroids the latter is as hard as the former.
@@
For binary matroids, there is no gap in approximability between the known-matroid MSP and the online-revealed-matroid MSP.
As our second result, we use the existence of OMEs into a 2-transitive host matroid as a tool to establish the approximate equivalence of known-matroid MSP and prophet MSP with pairwise-independent distributions.
Theorem 5.3.
Suppose a class $\mathcal{C}$ of matroids admits a randomized order-independent online matroid embedding into matroid BigM which is $2$-transitive. Then an $\alpha$-approximation to the prophet MSP with pairwise-independent weight distributions for $\textbf{BigM}_{[k]}$, implies that for some constant $C>0$ there is a $C\cdot(\alpha-o(1))$-approximation to the known-matroid MSP for $\mathcal{C}$.
Noting that the full binary matroid is $2$-transitive, and there exists an order-independent OME from the class of binary matroids into a full binary matroid, we obtain the following corollary.
Corollary 5.4.
For binary matroids, there is a constant-factor gap in approximability between the known-matroid MSP and the prophet MSP with pairwise-independent weight distributions.
We note that we can also chain the two reductions, and this way relate the online-revealed-matroid MSP to the Prophet MSP with pairwise-independent distributions. The rest of this section is devoted to the proofs of the reductions.
@@
Observe that if we ignore the weights, the order of the elements of BigM is the same in both processes since they are determined by the timestamps $t_{vj}$. What we are left to argue is that for each $v$ we select uniformly random timestamp $t_{vj}$ to assign the non-negative weight. For that, observe that since $f_{\textbf{M},\pi}$ is order independent, it has the same distribution as if we first sampled a monomorphism $f:\textbf{M}\rightarrow\textbf{BigM}$, and then we sampled an independent uniform indexing $\pi$ for the arrival order of the elements in $f(\textbf{M})$. This implies that when we assign the timestamps $T_{1},\dots,T_{n}$ according to $\pi$, the resulting distribution is the same as if we assigned i.i.d. Uniform[0,1][0,1] timestamps $T_{v}$ to each element $v\in f(\textbf{M})$, and therefore, the interval $T_{v}$ lands in is uniformly chosen and independent across elements $v\in f(\textbf{M})$. Now fix a certain timestamp $t_{vj}$ and let $I_{i}$ be the interval containing it. We will show that the probability that this timestamp is selected is exactly $1/k$.
Consider two cases: either $X_{vi}\leq k/d$ in which case the probability of sampling $t_{vj}$ is the probability that the timestamp $T_{v}$ is in $I_{i}$ (which is $1/d$), times the probability we decide to sample a timestamp inside $I_{i}$ (which is $X_{vi}d/k$), times the probability that out of those, we choose $t_{vj}$ (which is $1/X_{vi}$). The total probability is:
@@
Taking those lemmas together, we can conclude the proof of Theorem 5.1. For that, let Alg be an $\alpha$-competitive algorithm for $\textbf{BigM}_{[k]}$ and let $Y$ represent the sequence of the MSP sampled by the offline reduction. Let’s represent by $\textsf{Alg}(Y)$ the weight of the elements selected by Alg and Opt the weight of the optimal basis. By the fact that the offline reduction produces an instance in random order, we know that $\mathbb{E}[\textsf{Alg}(?^{\prime}Y)]\geq\alpha\textsf{Opt}$.
@@
5.2 Proof of Theorem 5.3
Let M be the unknown matroid with $|\textbf{M}|=n$ that is revealed to the algorithm in an online fashion, which admits an order-independent randomized OME into BigM with $|\textbf{BigM}|=M$. First, we obtain the following simple reduction that allows us to focus on a special class of the prophet MSP in which each element takes a weight from the set of weights $W$ with $|W|=m=O(n^{2})$. In addition, we can restrict the weight distribution such that each element has a distinct weight from the weight class.
Lemma 5.8.
If there exists an $\alpha$-approximation to the prophet MSP on M with weight distribution $\mathcal{D}$ supported over the set of weights $W$ with $\textsf{rank}(\textbf{M})=d$, $|\textbf{M}|=n$, $|W|=O(n^{2})$ with $\max_{w\in W}w\leq 1$ and $\mathbb{E}_{\mathcal{D}}[\textbf{OPT}(\textbf{M})]\in\left[\frac{1}{16},1\right]$ such that for all $w\in W$, there exists at most one element assigned weight of $w$ with probability one, then there exists an $\left(\frac{\alpha}{256}-\frac{1}{32d}\right)$-approximation to prophet MSP on matroid M with any arbitrary weight distribution.
The proof of the above lemma simply follows from Sublemma-4.2 from [19] that reduces any arbitrary prophet MSP with $O(\log(|\textsf{rank}(\textbf{M})|))$ many weights and $\mathbb{E}_{\mathcal{D}}[\textbf{OPT}(\textbf{M})]\in\left[\frac{1}{16},1\right]$. We then add distinct noise of the order of $O\left(\frac{1}{n^{2}}\right)$ to ensure that the weight of each element is distinct. The full proof of the reduction is delegated to Section A.
For simplicity, we let $W=\{w_{1},\dots,w_{m}\}$ and consider the prophet MSP on matroid M and weight distribution $\mathcal{D}$ supported over the set of weights $W$ satisfying the conditions from Lemma 5.8.
Extending BigM with Copies
We let $\textbf{BigM}_{[m\cdot N]}$ be a matroid with $m\cdot N$ parallel copies of each element of BigM with $|\textbf{BigM}|=M$ and integer $N=\Omega\left(2^{M^{2}}\right)$. We divide the set of $N\cdot m$ copies into $m$ sets of size $N$, each part corresponding to weight class $w_{i}$. We use $N_{i}$ to denote the set of labels corresponding to weight $w_{i}$ for all $i\in[m]$ with $|N_{i}|=N$. We sometimes denote $N_{i}$ by $N]=\{1,2,\dots,N\}$ whenever it is clear from the context which $w_{i}$ we are referring to. In addition, the $\ell$-th copy of the weight class corresponding to weight $w_{i}$ of element $\mathbf{v}\in\textbf{BigM}$ is denoted as $\mathbf{v}^{i,\ell}$.
Reduction to “Almost” Pairwise Independent Prophet MSP
We first define the weight distribution $\mathcal{D}^{*}$ over $\textbf{BigM}_{[m\cdot N]}$ in Definition 1, which is “almost” pairwise independent. Then in Theorem 5.9, we show an existence of exact pairwise independent weight distribution $\tilde{\mathcal{D}}$ over $\textbf{BigM}_{[m\cdot N]}$ which is “close” to the distribution defined in Definition 1 in total-variation distance. This allows us to utilize the fact that any algorithm $\mathcal{A}$ can not distinguish between the almost pairwise independent weight distribution $\mathcal{D}^{*}$ and $\tilde{\mathcal{D}}$ with high probability. Finally, we complete the proof of Theorem 5.3.
We begin by defining the almost pairwise independent weight distribution over $\textbf{BigM}_{[m\cdot N]}$.
Definition 1 (Almost P.W. Independent Distribution).
Consider the weight distribution $\mathcal{D}^{*}$ over the elements of $\textbf{BigM}_{[m\cdot N]}$ defined as follows:
  1. 1.
    Given an order independent OMM $f^{\prime}:\textbf{M}\rightarrow\textbf{BigM}$, we sample a random automorphism $f^{\prime\prime}\in\operatorname{Aut}(\textbf{BigM})$ and obtain an order independent matroid morphism $f=f^{\prime\prime}\circ f^{\prime}$.
  2. 2.
    For any $\mathbf{v}\in\textbf{M}$ with $w(\mathbf{v})=w_{i}$, let $\mathbf{u}=f(\mathbf{v})$. We sample $\ell\sim\operatorname{Unif}(N_{i})$ and assign the weight of $w(\mathbf{u}^{i,\ell})=w_{i}$.
  3. 3.
    We assign the weight of the rest of the elements of $\textbf{BigM}_{[m\cdot N]}$ to be zero.
We first observe that when $N$ is much larger than $M$, the distribution in the above definition is almost pairwise independent. To see this, we first observe that for any $i\in[m],\ell\in N_{i}$ and $\mathbf{u}\in\textbf{BigM}$, $\mathbf{u}^{i,\ell}\in\textbf{BigM}_{[m\cdot N]}$ can potentially either take a weight of $w_{i}$ or zero. For simplicity, now consider any two distinct elements $\mathbf{u}_{i,\ell},\mathbf{u}^{\prime}_{j,\ell^{\prime}}\in\textbf{BigM}_{[m\cdot N]}$ and weight distribution $\mathcal{D}$ such that there always exists a pair of elements $\mathbf{v},\mathbf{v}^{\prime\prime}\in\textbf{M}$ that are assigned weights of $w_{i},w_{j}$, respectively. Since $f^{\prime}$ is a random automorphism $\operatorname{Aut}(\textbf{BigM})$, we have $\Pr[w(\mathbf{u}_{i,\ell})=w_{i}]=\Pr[f(\mathbf{v})=\mathbf{u}]\cdot\frac{1}{N}=\frac{1}{M\cdot N}$. On the other hand, we have
$\Pr[w(\mathbf{u}_{i,\ell})=w_{i}\land w(\mathbf{u}_{j,\ell})=w_{j}]=\Pr[f^{\prime}(\mathbf{v})=\mathbf{v}\land f^{\prime}(\mathbf{v}^{\prime})=\mathbf{u}^{\prime}]\cdot\frac{1}{N^{2}}=\frac{1}{M(M-1)}\cdot\frac{1}{N^{2}},$
which is close to the product $\Pr[w(\mathbf{u}_{i,\ell})=w_{i}]\cdot\Pr[w(\mathbf{u}_{j,\ell})=w_{j}]=\frac{1}{M^{2}\cdot N^{2}}$. However, in general, we can not guarantee that $\mathcal{D}$ will always assign weights $w_{i},w_{j}$ to some pair of elements of $\mathbf{v},\mathbf{v}^{\prime}$ of M. In addition, the above argument also fails if we have $i=j$ or $\mathbf{u}=\mathbf{u}^{\prime}$ as in both of these cases, $\Pr[w(\mathbf{u}_{i,\ell})=w_{i}\land w(\mathbf{u}_{j,\ell})=w_{j}]=0$. Intuitively, we can circumvent these pairwise correlation issues by taking $N$ large enough as it makes pairwise correlations small enough. More precisely,
$|\Pr[w(\mathbf{u}_{i,\ell})=w_{i}\land w(\mathbf{u}_{j,\ell})=w_{j}]-\Pr[w(\mathbf{u}_{i,\ell})=w_{i}]\cdot\Pr[w(\mathbf{u}_{j,\ell})=w_{j}]|=O\left(\frac{1}{M^{2}\cdot N^{2}}\right).$
Using this observation, we prove the following technical theorem.
Theorem 5.9.
For $N=\Omega\left(2^{M^{2}}\right)$, let weight distribution $\mathcal{D}^{*}$ over $\textbf{BigM}_{[m\cdot N]}$ be defined as in Definition 1, then there exist a pairwise-independent weight distribution $\tilde{\mathcal{D}}$ over $\textbf{BigM}_{[m\cdot N]}$ such that $\operatorname{TV}_{\mathcal{D}^{*},\tilde{\mathcal{D}}}\leq O\left(\frac{m^{3}}{M}\right)$.
The proof of the above theorem is highly technical and constructs an explicit $\tilde{\mathcal{D}}$ by a sequence of small perturbations to $\mathcal{D}^{*}$. For the sake of the uninterrupted flow of the presentation, we delegate it to Appendix B. We emphasize that the choice of $N=\Omega(2^{M^{2}})$ is required due to the limitations of the techniques developed to prove Theorem 5.9. We conjecture that one can prove the similar theorem for $N=\Omega(\operatorname{Poly}(M))$, which we leave as an intriguing technical open problem.
Proof of Theorem 5.3
We now complete the proof of the main theorem.
Proof of Theorem 5.3.
To prove the main theorem, we first prove the following: if there exists an $\alpha$-approximate algorithm to the prophet MSP instance with a pairwise independent weight distribution for the matroid $\textbf{BigM}_{[m\cdot N]}$, then there exists an $(\alpha-o(1))$-approximate algorithm for the prophet MSP instance for the matroid M with weight distribution $\mathcal{D}$ supported over the set of weights $W$ with $\textsf{rank}(\textbf{M})=d$, $|\textbf{M}|=n$ and $|W|=O(n^{2})$ with $\max_{w\in W}w\leq 1$ and $\mathbb{E}_{\mathcal{D}}[\textbf{OPT}(\textbf{M})]\in\left[\frac{1}{16},1\right]$ such that for all $w\in W$, there exists at most one element assigned weight of $w$ with probability one. Combining this with Lemma 5.8, we will conclude the proof of the Theorem 5.3
Given an order-independent matroid morphism $f^{\prime}:\textbf{M}\rightarrow\textbf{BigM}$, we sample a random automorphism $f^{\prime\prime}\in\operatorname{Aut}(\textbf{BigM})$ and obtain an order-independent matroid morphism $\tilde{f}=f^{\prime\prime}\circ f^{\prime}$. Given $\tilde{f}:\textbf{M}\rightarrow\textbf{BigM}$, we obtain $f:\textbf{M}\rightarrow\textbf{BigM}_{[m]}$ that maps each $\mathbf{v}\in\textbf{M}$ to $\mathbf{u}^{i}\in\textbf{BigM}_{[m]}$ iff $\tilde{f}(\mathbf{v})=\mathbf{u}$ and $w(\mathbf{v})=w_{i}$. We then consider $\textbf{BigM}_{[m\cdot N]}$, i.e. matroid $\textbf{BigM}_{[m]}$ with $N$ many copies of each element.
We first consider an offline reduction as follows: for all $\mathbf{v}^{i}\in\textbf{BigM}_{[m]}$, sample $N$ many independent arrival times from $\operatorname{Unif}[0,1]$ denoting the uniformly random arrival times of elements of $\textbf{BigM}_{[m\cdot N]}$. Given any $\mathbf{v}\in\textbf{M}$ and pair $(\mathbf{v},w(\mathbf{v})=w_{i})$, with $\mathbf{u}^{i}=f(\mathbf{v})$, let $w(\mathbf{u}^{i,\ell})=w_{i}$ uniformly random from $\ell\in[N]$.
Since $f$ does not require any information about the arrival order, the above-described offline reduction is a valid instance of matroid prophet secretary over $\textbf{BigM}_{[m\cdot N]}$. In addition, the induced weight assignment over BigM due to offline reduction is identical to the distribution $\mathcal{D}^{*}$ defined in Definition 1.
Given the uniformly random arrival of elements of M, we construct an “almost online implementation” of the above offline reduction similar to the proof of Theorem 5.1. We let $N$ be large enough ($\Omega(n^{2}/\varepsilon)$) such that the probability of “almost online implementation” raising a flag is at most $\varepsilon$. In fact, for the proof of Theorem 5.9, we let $N=\Omega(2^{M^{2}})$ which satisfies the required condition.
We let $\mathcal{A}$ be an $\alpha$-approximate algorithm for the pairwise-independent prophet MSP on matroid $\textbf{BigM}_{[m\cdot N]}$. Since there exists a pairwise independent weight distribution $\tilde{\mathcal{D}}$ over BigM within the total variation distance of $O\left(\frac{m^{3}}{M}\right)$, the algorithm $\mathcal{A}$ can not distinguish the weight distribution $\mathcal{D}^{*}$ from $\tilde{\mathcal{D}}$ with probability at least $1-O\left(\frac{m^{2}}{M}\right)$.
Let $\mathcal{E}$ be the event when the algorithm $\mathcal{A}$ can not distinguish between $\mathcal{D}^{*}$ and $\tilde{D}$. We note that when event $\mathcal{E}$ does not hold, the offline optimal can be bounded by
$\mathbb{E}[\textbf{OPT}(\textbf{BigM}_{m\cdot N})\mid\mathcal{E}^{c}]\leq\textsf{rank}(\textbf{BigM}_{[m\cdot N]})\cdot\max_{w_{i}\in W}{w_{i}}\leq n\cdot 1=n,$
where the second inequality follows because $\textsf{rank}(\textbf{BigM}_{[m\cdot N]})=n$ and $w_{i}\leq 1$ for all $i\in[m\cdot N]$.
Now, let $S$ be the selected set of elements of BigM by $\mathcal{A}$ w.r.t. weight distribution $\mathcal{D}^{*}$. We can bound,
$\mathbb{E}[w(S)]$ $=\mathbb{E}[w(S)\mid\mathcal{E}]\cdot\Pr[\mathcal{E}]+\mathbb{E}[w(S)\mid\mathcal{E}^{c}]\cdot\Pr[\mathcal{E}^{c}]$
$\leq\mathbb{E}[w(S)\mid\mathcal{E}]+\mathbb{E}[\textbf{OPT}(\textbf{BigM}_{[M\cdot m]})\mid\mathcal{E}^{c}]\cdot\frac{m^{3}}{M}$
$\leq\mathbb{E}[w(S)\mid\mathcal{E}]+\frac{m^{3}\cdot n}{M}.$
Above, the first inequality holds because $\Pr[\mathcal{E}^{c}]\leq\frac{m^{2}}{M}$ and the second inequality holds because $\mathbb{E}[\textbf{OPT}(\textbf{BigM}_{[m\cdot N]})\mid\mathcal{E}^{c}]\leq n$. Due to our reduction, the performance of the algorithm on the original Prophet MSP instance $\mathbb{F}^{\prime}$ is lower bounded by $\mathbb{E}[w(S)\mid\mathcal{E}]$, next we lower bound the expectation $\mathbb{E}[w(S)\mid\mathcal{E}]$,
$\mathbb{E}[w(S)\mid\mathcal{E}]$ $\geq\mathbb{E}[w(S)]-\frac{m^{3}\cdot n}{M}$
$\geq\alpha\cdot\mathbb{E}_{\tilde{\mathcal{D}}}[\textbf{OPT}(\textbf{BigM}_{[m\cdot N]})]-o(1)\cdot\textbf{OPT}(\textbf{M})$
$\geq\textbf{OPT}(\textbf{M})\cdot\left(\alpha-\frac{\alpha\cdot m^{3}}{M}\right)-o(1)\cdot\textbf{OPT}(\textbf{M})\geq\left(\alpha-o(1)\right)\cdot\textbf{OPT}(\textbf{M}).$
Above, the second inequality holds because $\mathcal{A}$ is an $\alpha$ approximate algorithm for Prophet MSP with pairwise independent prior and $\frac{m^{3}\cdot n}{M}=O(m^{4}/2^{m})=o(1)$. The third inequality holds because $\operatorname{TV}_{\mathcal{D}^{*},\tilde{\mathcal{D}}}\leq O(m^{3}/M)$. Finally, since the performance of the reduction on the original matroid secretary is
$\mathbb{E}[w(S)\mid\mathcal{E}]\cdot\Pr[\text{Reduction does not Flag}]\geq(1-\varepsilon)\cdot\left(\alpha-o(1)\right)\cdot\textbf{OPT}(\textbf{M}).$
This concludes the proof. ∎
6 OMEs Beyond Binary Matroids
We next explore OMEs beyond binary matroids. In Section 6.1, we show that there cannot be an OME that embeds graphical matroids into graphical matroids or, more generally, regular matroids. Then, in Section 6.2, we give an OME that embeds laminar matroids into laminar matroids. Finally, in Section 6.3, we show that there is no universal host matroid, that allows embedding of all matroids on $n$ elements of a given rank.
6.1 Graphic and Regular Matroids